# CADmium: Fine-Tuning Code Language Models for Text-Driven Sequential CAD Design

### v1

@@ 5 Discussion and Conclusion
Despite the current progress of foundation models that have multimodal capabilities, achieving state-of-the-art in language, image, and 3D object generation [31] and understanding, there is still plenty of room for improvement when it comes to domain-specific tasks. Much of this improvement depends on the availability of domain-relevant data and the capacity of large machine learning models to learn from it. Fine-tuning large language models for generating application-specific 3D structural entities, such as small molecules, proteins, and materials, is an active area of research that has shown remarkable progress in recent years. Likewise, CAD also deals with 3D mechanical objects, and automating this process can rapidly advance manufacturing sectors and 3D printing. Specifically, text-conditioned CAD model generation and completion can greatly help engineers and designers to quickly optimize the design process with simple natural language instructions.
With the absence of large and good-quality human-annotated textual descriptions of CAD objects that are publicly available, it is necessary to utilize powerful and multimodal language models that can generate human-level annotations given the images and design history of CAD objects. In this regard our main goals were to further the state-of-the-art in terms of the quality of the text annotations, improve the training process by using instruction-tuned Code Qwen models, and augment the evaluation pipeline by including additional metrics that can provide deeper structural and topological insights (Appendix D). We also highlight some important limitations in our work. Although we demonstrated that GPT-generated prompts exhibit human-like characteristics, further investigation is needed to evaluate their generalization to diverse types of human prompts. Additionally, we observed a few outlier reconstructions with high Chamfer distances from the ground truths, which may be mitigated by increasing the training data and model size. In future work, we plan to extend our approach with a multi-modal framework that facilitates editing CAD designs.
@@ Appendix A Example Annotations
This appendix provides illustrative examples comparing our CADmium annotations, generated using GPT-4.1, with the expert-level annotations from the Text2CAD dataset for the same CAD models. These qualitative comparisons, presented in Figures 3 through 7, highlight several improvements discussed in Section 3.1 and quantitatively supported by the analysis in Figure 2.
@@ Appendix B Metrics on a Common Evaluable Subset
Table 5: Performance comparison on the human-annotated CADPrompt dataset (common subset evaluation). Metrics are computed on the intersection of samples for which both the models, produced valid CAD sequences.
Model Text2CAD Qwen Coder 14B
@@
Line F1 $\uparrow$ 0.67 $\pm$ 0.27 0.69 $\pm$ 0.31
Arc F1 $\uparrow$ 0.00 $\pm$ 0.00 0.23 $\pm$ 0.33
Circle F1 $\uparrow$ 0.33 $\pm$ 0.40 0.61 $\pm$ 0.43
Extrusion F1 $\uparrow$ 0.84 $\pm$ 0.20 0.89 $\pm$ 0.16
CD median $\downarrow$ 129.47 112.84
SIR $\downarrow$ 0.03 $\pm$ 0.13 0.10 $\pm$ 0.27
DangEL $\downarrow$ 0.21 $\pm$ 1.52 2.99 $\pm$ 10.09
SegE $\downarrow$ 0.10 $\pm$ 0.51 0.78 $\pm$ 2.27
FluxEE (×$\times 10^{2}$) $\downarrow$ 0.05 $\pm$ 0.44 2.28 $\pm$ 9.26
Watertightness (%) $\uparrow$ 95.74 83.80
@@
EECM $\uparrow$ 0.66 0.67
DMCD (×$\times 10^{3}$) $\downarrow$ 14.88 $\pm$ 12.93 8.34 $\pm$ 11.00
SD mean (×$\times 10^{2}$) $\downarrow$ 15.69 $\pm$ 14.47 7.22 $\pm$ 12.78
SD median (×$\times 10^{2}$) $\downarrow$ 12.26 1.51
Table 6: Impact of Qwen Coder model scale (1.5B, 3B, 7B, 14B parameters) on the CADmium test set (common subset evaluation). For each dataset, metrics are computed on the intersection of samples for which all three model sizes produced valid CAD sequences. All models were trained on CADmium annotations.
@@
Line F1 $\uparrow$ 0.93 $\pm$ 0.17 0.94 $\pm$ 0.17 0.94 $\pm$ 0.16 0.95 $\pm$ 0.14
Arc F1 $\uparrow$ 0.78 $\pm$ 0.34 0.80 $\pm$ 0.35 0.82 $\pm$ 0.34 0.86 $\pm$ 0.30
Circle F1 $\uparrow$ 0.90 $\pm$ 0.24 0.93 $\pm$ 0.20 0.93 $\pm$ 0.21 0.95 $\pm$ 0.17
Extrusion F1 $\uparrow$ 0.99 $\pm$ 0.06 0.99 $\pm$ 0.06 0.99 $\pm$ 0.05 0.99 $\pm$ 0.05
CD median $\downarrow$ 0.26 0.21 0.23 0.20
SIR $\downarrow$ 0.04 $\pm$ 0.15 0.03 $\pm$ 0.14 0.03 $\pm$ 0.13 0.02 $\pm$ 0.13
DangEL $\downarrow$ 0.93 $\pm$ 5.83 0.76 $\pm$ 4.40 0.65 $\pm$ 4.15 0.57 $\pm$ 3.81
SegE $\downarrow$ 0.33 $\pm$ 1.71 0.30 $\pm$ 1.61 0.26 $\pm$ 1.13 0.24 $\pm$ 1.33
FluxEE (×$\times 10^{2}$) $\downarrow$ 0.30 $\pm$ 3.20 0.30 $\pm$ 3.47 0.24 $\pm$ 2.94 0.25 $\pm$ 2.84
Watertightness (%) $\uparrow$ 90.38 90.31 90.54 90.78
Num Watertight 5433 5433 5433 5433
EECM $\uparrow$ 0.90 0.91 0.91 0.92
DMCD (×$\times 10^{3}$) $\downarrow$ 1.98 $\pm$ 6.00 1.75 $\pm$ 6.05 1.69 $\pm$ 5.98 1.55 $\pm$ 5.79
SD mean (×$\times 10^{2}$) $\downarrow$ 2.12 $\pm$ 7.51 1.58 $\pm$ 6.12 1.56 $\pm$ 6.27 1.36 $\pm$ 5.83
SD median (×$\times 10^{2}$) $\downarrow$ 0.00 0.00 0.00 0.00
Table 7: Impact of Qwen Coder model scale (1.5B, 3B, 7B, 14B parameters) on the Fusion360 test set (common subset evaluation). For each dataset, metrics are computed on the intersection of samples for which all three model sizes produced valid CAD sequences. All models were trained on CADmium annotations.
@@
Line F1 $\uparrow$ 0.86 $\pm$ 0.25 0.88 $\pm$ 0.24 0.90 $\pm$ 0.22 0.91 $\pm$ 0.22
Arc F1 $\uparrow$ 0.74 $\pm$ 0.35 0.75 $\pm$ 0.37 0.78 $\pm$ 0.35 0.81 $\pm$ 0.33
Circle F1 $\uparrow$ 0.90 $\pm$ 0.23 0.92 $\pm$ 0.19 0.92 $\pm$ 0.17 0.94 $\pm$ 0.16
Extrusion F1 $\uparrow$ 0.98 $\pm$ 0.06 0.98 $\pm$ 0.07 0.99 $\pm$ 0.05 0.99 $\pm$ 0.05
CD median $\downarrow$ 103.82 72.37 85.95 73.71
SIR $\downarrow$ 0.04 $\pm$ 0.15 0.04 $\pm$ 0.15 0.05 $\pm$ 0.17 0.03 $\pm$ 0.13
DangEL $\downarrow$ 1.02 $\pm$ 4.78 0.85 $\pm$ 4.47 1.21 $\pm$ 5.67 0.73 $\pm$ 4.35
SegE $\downarrow$ 0.37 $\pm$ 1.30 0.34 $\pm$ 1.43 0.38 $\pm$ 1.42 0.28 $\pm$ 1.26
FluxEE (×$\times 10^{2}$) $\downarrow$ 0.20 $\pm$ 2.22 0.20 $\pm$ 2.33 0.18 $\pm$ 2.37 0.18 $\pm$ 2.01
Watertightness (%) $\uparrow$ 86.62 86.50 85.21 87.35
Num Watertight 3957 3957 3957 3957
EECM $\uparrow$ 0.82 0.84 0.84 0.87
DMCD (×$\times 10^{3}$) $\downarrow$ 3.25 $\pm$ 7.76 2.83 $\pm$ 7.41 2.80 $\pm$ 7.12 2.70 $\pm$ 7.78
SD mean (×$\times 10^{2}$) $\downarrow$ 3.02 $\pm$ 8.83 2.14 $\pm$ 6.77 2.61 $\pm$ 8.12 2.19 $\pm$ 7.23
SD median (×$\times 10^{2}$) $\downarrow$ 0.00 0.00 0.00 0.00
Table 8: Impact of annotation quality using the Text2CAD architecture (common subset evaluation). Models were trained on either CADmium or Text2CAD annotations. For each of the four training/evaluation set pairings, metrics are computed on the intersection of samples for which both compared conditions produced valid CAD sequences.
Trained on CADmium Text2CAD
Evaluated on CADmium Text2CAD CADmium Text2CAD
@@
Line F1 $\uparrow$ 0.89 $\pm$ 0.21 0.74 $\pm$ 0.29 0.64 $\pm$ 0.25 0.83 $\pm$ 0.26
Arc F1 $\uparrow$ 0.60 $\pm$ 0.41 0.22 $\pm$ 0.37 0.25 $\pm$ 0.35 0.38 $\pm$ 0.43
Circle F1 $\uparrow$ 0.89 $\pm$ 0.26 0.58 $\pm$ 0.44 0.47 $\pm$ 0.39 0.76 $\pm$ 0.37
Extrusion F1 $\uparrow$ 0.98 $\pm$ 0.07 0.90 $\pm$ 0.18 0.81 $\pm$ 0.17 0.94 $\pm$ 0.15
CD median $\downarrow$ 0.31 50.78 33.38 0.37
SIR $\downarrow$ 0.04 $\pm$ 0.16 0.04 $\pm$ 0.17 0.08 $\pm$ 0.22 0.04 $\pm$ 0.17
DangEL $\downarrow$ 1.36 $\pm$ 6.45 1.11 $\pm$ 6.32 2.38 $\pm$ 9.17 1.06 $\pm$ 5.59
SegE $\downarrow$ 0.63 $\pm$ 5.61 0.50 $\pm$ 2.85 1.06 $\pm$ 4.24 0.48 $\pm$ 2.08
FluxEE (×$\times 10^{2}$) $\downarrow$ 0.39 $\pm$ 3.49 0.57 $\pm$ 5.78 0.93 $\pm$ 5.84 0.42 $\pm$ 4.41
Watertightness (%) $\uparrow$ 87.63 91.90 82.48 90.52
@@
EECM $\uparrow$ 0.87 0.77 0.66 0.82
DMCD (×$\times 10^{3}$) $\downarrow$ 3.37 $\pm$ 8.55 7.64 $\pm$ 12.59 9.19 $\pm$ 11.97 5.09 $\pm$ 10.70
SD mean (×$\times 10^{2}$) $\downarrow$ 2.50 $\pm$ 6.63 8.31 $\pm$ 12.08 11.31 $\pm$ 15.16 4.68 $\pm$ 10.00
SD median (×$\times 10^{2}$) $\downarrow$ 0.00 3.20 5.07 0.02

### v2

@@ 5 Discussion and Conclusion
Despite the current progress of foundation models that have multimodal capabilities, achieving state-of-the-art in language, image, and 3D object generation (Xiang et al., 2024) and understanding, there is still plenty of room for improvement when it comes to domain-specific tasks. Much of this improvement depends on the availability of domain-relevant data and the capacity of large machine learning models to learn from it. Fine-tuning large language models for generating application-specific 3D structural entities, such as small molecules, proteins, and materials, is an active area of research that has shown remarkable progress in recent years. Likewise, CAD also deals with 3D mechanical objects, and automating this process can rapidly advance manufacturing sectors and 3D printing. Specifically, text-conditioned CAD model generation and completion can greatly help engineers and designers to quickly optimize the design process with simple natural language instructions.
With the absence of large and good-quality human-annotated textual descriptions of CAD objects that are publicly available, it is necessary to utilize powerful and multimodal language models that can generate human-level annotations given the images and design history of CAD objects. In this regard our main goals were to further the state-of-the-art in terms of the quality of the text annotations, improve the training process by using instruction-tuned Code Qwen models, and augment the evaluation pipeline by including additional metrics that can provide deeper structural and topological insights (Appendix D). We also highlight some important limitations in our work. Although we demonstrated that GPT-generated prompts exhibit human-like characteristics, further investigation is needed to evaluate their generalization to diverse types of human prompts. Additionally, we observed a few outlier reconstructions with high Chamfer distances from the ground truths, which may be mitigated by increasing the training data and model size. In future work, we plan to extend our approach with a multimodal framework that facilitates editing CAD designs.
@@ 6 Acknowledgments
Figure 3: Examples of generated objects from the CADmium test set: 3D models are generated by both Qwen2.5 Coder (fine-tuned) and Text2CAD (trained) on the CADmium training set. The examples are shown using CADmium’s expert prompts, with the natural-language descriptions shortened for readability. We present both successful and failed reconstructions from Qwen2.5 Coder and compare them against the outputs of Text2CAD.
@@ Appendix A Example Annotations
This appendix provides illustrative examples comparing our CADmium annotations, generated using GPT-4.1, with the expert-level annotations from the Text2CAD dataset for the same CAD models. These qualitative comparisons, presented in Figures 4 through 8, highlight several improvements discussed in Section 3.1 and quantitatively supported by the analysis in Figure 2.
@@ Appendix B Metrics on a Common Evaluable Subset
Table 5: Performance comparison on the human-annotated CADPrompt dataset (common subset evaluation). Metrics are computed on the intersection of samples for which both models produced valid CAD sequences. Reported values include 95% confidence intervals in brackets. Boldface indicates the better-performing model whenever the difference is statistically significant at $p<0.05$ (Wilcoxon signed-rank test for continuous metrics, McNemar’s test for the binary watertightness metric).
Model Text2CAD Qwen-2.5 Coder 14B
@@
Line F1 $\uparrow$ 0.67 [$0.63$, $0.72$] 0.69 [$0.64$, $0.74$]
Arc F1 $\uparrow$ 0.00 [$0.00$, $0.00$] 0.23 [$0.11$, $0.36$]
Circle F1 $\uparrow$ 0.33 [$0.23$, $0.42$] 0.61 [$0.51$, $0.71$]
Extrusion F1 $\uparrow$ 0.84 [$0.82$, $0.87$] 0.89 [$0.87$, $0.92$]
CD mean $\downarrow$ 224.38 [$187.33$, $263.47$] 203.04 [$173.25$, $235.10$]
CD median $\downarrow$ 129.47 [$87.59$, $170.25$] 112.84 [$78.86$, $177.11$]
SIR $\downarrow$ 0.03 [$0.01$, $0.05$] 0.10 [$0.07$, $0.15$]
DangEL $\downarrow$ 0.21 [$0.03$, $0.47$] 2.99 [$1.67$, $4.64$]
SegE $\downarrow$ 0.10 [$0.03$, $0.18$] 0.78 [$0.48$, $1.14$]
FluxEE ($\times 10^{2}$) $\downarrow$ 0.05 [$0.00$, $0.12$] 2.28 [$1.04$, $3.78$]
Watertightness (%) $\uparrow$ 95.74 [$92.55$, $98.40$] 83.80 [$78.21$, $88.83$]
@@
EECM $\uparrow$ 0.66 [$0.55$, $0.71$] [$0.59$, $0.75$]
DMCD ($\times 10^{3}$) $\downarrow$ 14.88 [$12.74$, $17.12$] 8.34 [$6.54$, $10.32$]
SD mean ($\times 10^{2}$) $\downarrow$ 15.69 [$13.37$, $18.27$] 7.22 [$5.09$, $9.60$]
SD median ($\times 10^{2}$) $\downarrow$ 12.26 [$9.26$, $15.14$] 1.51 [$1.06$, $2.15$]
Table 6: Impact of Qwen Coder model scale (1.5B, 3B, 7B, 14B parameters) on the CADmium test set (common subset evaluation). For each dataset, metrics are computed on the intersection of samples for which all three model sizes produced valid CAD sequences. All models were trained on CADmium annotations. Reported values include 95% confidence intervals in brackets. Statistical significance is assessed with statistical tests on adjacent sizes (1.5B→3B, 3B→7B, 7B→14B). Bold indicates a size that is significantly better than the next smaller size at $p<0.05$. Details of the statistical testing procedure are provided in Section 4.
@@
Line F1 $\uparrow$ 0.93 [$0.92$, $0.93$] 0.94 [$0.93$, $0.94$] 0.94 [$0.94$, $0.95$] 0.95 [$0.95$, $0.95$]
Arc F1 $\uparrow$ 0.78 [$0.76$, $0.81$] 0.80 [$0.77$, $0.82$] 0.82 [$0.79$, $0.84$] 0.86 [$0.84$, $0.88$]
Circle F1 $\uparrow$ 0.90 [$0.89$, $0.91$] 0.93 [$0.92$, $0.94$] 0.93 [$0.92$, $0.94$] 0.95 [$0.94$, $0.95$]
Extrusion F1 $\uparrow$ 0.99 [$0.98$, $0.99$] 0.99 [$0.99$, $0.99$] 0.99 [$0.99$, $0.99$] 0.99 [$0.99$, $0.99$]
CD mean $\downarrow$ 64.74 [$60.84$, $68.67$] 39.48 [$36.03$, $42.97$] 54.23 [$50.43$, $58.22$] 37.11 [$34.13$, $40.33$]
CD median $\downarrow$ 0.26 [$0.25$, $0.27$] 0.21 [$0.21$, $0.22$] 0.23 [$0.21$, $0.24$] 0.20 [$0.20$, $0.21$]
SIR $\downarrow$ 0.04 [$0.03$, $0.04$] 0.03 [$0.03$, $0.03$] 0.03 [$0.02$, $0.03$] 0.02 [$0.02$, $0.03$]
DangEL $\downarrow$ 0.94 [$0.80$, $1.09$] 0.76 [$0.65$, $0.87$] 0.65 [$0.56$, $0.76$] 0.57 [$0.48$, $0.67$]
SegE $\downarrow$ 0.33 [$0.29$, $0.38$] 0.30 [$0.26$, $0.34$] 0.26 [$0.23$, $0.29$] 0.24 [$0.21$, $0.28$]
FluxEE ($\times 10^{2}$) $\downarrow$ 0.30 [$0.23$, $0.38$] 0.30 [$0.22$, $0.38$] 0.24 [$0.17$, $0.32$] 0.25 [$0.18$, $0.32$]
Watertightness (%) $\uparrow$ 90.38 [$89.69$, $91.06$] 90.31 [$89.62$, $90.99$] 90.54 [$89.87$, $91.19$] 90.78 [$90.13$, $91.44$]
Num Watertight 6205 6408 6575 6779
EECM $\uparrow$ 0.90 [$0.89$, $0.91$] 0.91 [$0.91$, $0.92$] 0.91 [$0.90$, $0.92$] 0.92 [$0.92$, $0.93$]
DMCD ($\times 10^{3}$) $\downarrow$ 2.42 [$2.25$, $2.60$] 2.18 [$2.00$, $2.37$] 2.08 [$1.91$, $2.26$] 1.95 [$1.78$, $2.14$]
SD mean ($\times 10^{2}$) $\downarrow$ 2.51 [$2.31$, $2.72$] 1.83 [$1.67$, $1.99$] 1.82 [$1.66$, $1.99$] 1.60 [$1.44$, $1.76$]
SD median ($\times 10^{2}$) $\downarrow$ 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$]
Table 7: Impact of Qwen Coder model scale (1.5B, 3B, 7B, 14B parameters) on the Fusion360 test set (common subset evaluation). For each dataset, metrics are computed on the intersection of samples for which all three model sizes produced valid CAD sequences. All models were trained on CADmium annotations. Reported values include 95% confidence intervals in brackets. Statistical significance is assessed with statistical tests on adjacent sizes (1.5B→3B, 3B→7B, 7B→14B). Bold indicates a size that is significantly better than the next smaller size at $p<0.05$. Details of the statistical testing procedure are provided in Section 4.
@@
Line F1 $\uparrow$ 0.86 [$0.85$, $0.87$] 0.88 [$0.87$, $0.89$] 0.90 [$0.89$, $0.91$] 0.91 [$0.90$, $0.92$]
Arc F1 $\uparrow$ 0.74 [$0.71$, $0.76$] 0.75 [$0.73$, $0.77$] 0.78 [$0.76$, $0.80$] 0.81 [$0.79$, $0.83$]
Circle F1 $\uparrow$ 0.90 [$0.89$, $0.91$] 0.92 [$0.92$, $0.93$] 0.92 [$0.92$, $0.93$] 0.94 [$0.93$, $0.94$]
Extrusion F1 $\uparrow$ 0.98 [$0.98$, $0.99$] 0.98 [$0.98$, $0.98$] 0.99 [$0.99$, $0.99$] 0.99 [$0.99$, $0.99$]
CD mean $\downarrow$ 227.21 [$218.41$, $236.09$] 188.94 [$181.02$, $197.01$] 206.18 [$197.73$, $214.94$] 211.93 [$202.79$, $221.16$]
CD median $\downarrow$ 103.82 [$97.32$, $110.25$] 72.37 [$64.80$, $80.78$] 85.95 [$79.99$, $91.91$] 73.71 [$63.45$, $82.08$]
SIR $\downarrow$ 0.04 [$0.03$, $0.04$] 0.04 [$0.03$, $0.04$] 0.05 [$0.04$, $0.05$] 0.03 [$0.03$, $0.03$]
DangEL $\downarrow$ 1.03 [$0.90$, $1.15$] 0.86 [$0.73$, $0.98$] 1.21 [$1.06$, $1.37$] 0.73 [$0.61$, $0.85$]
SegE $\downarrow$ 0.37 [$0.34$, $0.41$] 0.34 [$0.30$, $0.38$] 0.38 [$0.34$, $0.42$] 0.28 [$0.25$, $0.32$]
FluxEE ($\times 10^{2}$) $\downarrow$ 0.20 [$0.15$, $0.27$] 0.20 [$0.14$, $0.27$] 0.18 [$0.12$, $0.25$] 0.18 [$0.12$, $0.23$]
Watertightness (%) $\uparrow$ 86.62 [$85.78$, $87.46$] 86.50 [$85.67$, $87.35$] 85.21 [$84.38$, $86.04$] 87.35 [$86.56$, $88.10$]
Num Watertight 5241 5192 5504 5941
EECM $\uparrow$ 0.82 [$0.81$, $0.83$] 0.84 [$0.83$, $0.86$] 0.84 [$0.83$, $0.85$] 0.87 [$0.86$, $0.88$]
DMCD ($\times 10^{3}$) $\downarrow$ 4.04 [$3.78$, $4.30$] 3.36 [$3.12$, $3.61$] 3.56 [$3.31$, $3.81$] 3.26 [$3.02$, $3.52$]
SD mean ($\times 10^{2}$) $\downarrow$ 3.55 [$3.29$, $3.82$] 2.53 [$2.33$, $2.75$] 2.91 [$2.68$, $3.15$] 2.51 [$2.29$, $2.73$]
SD median ($\times 10^{2}$) $\downarrow$ 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$] 0.00 [$0.00$, $0.00$]
Table 8: Impact of annotation quality using the Text2CAD architecture (common subset evaluation). Models were trained on either CADmium or Text2CAD annotations. For each of the four training/evaluation set pairings, metrics are computed on the intersection of samples for which both compared conditions produced valid CAD sequences. Reported values include 95% confidence intervals in brackets. Boldface indicates statistically significant differences ($p<0.05$); details of the statistical testing procedure are provided in Section 4.
Evaluated on CADmium Text2CAD
Trained on CADmium Text2CAD CADmium Text2CAD
@@
Line F1 $\uparrow$ 0.89 [$0.89$, $0.90$] 0.74 [$0.73$, $0.75$] 0.64 [$0.64$, $0.65$] 0.83 [$0.82$, $0.83$]
Arc F1 $\uparrow$ 0.60 [$0.58$, $0.62$] 0.22 [$0.20$, $0.24$] 0.25 [$0.23$, $0.27$] 0.38 [$0.36$, $0.41$]
Circle F1 $\uparrow$ 0.89 [$0.88$, $0.90$] 0.58 [$0.57$, $0.60$] 0.47 [$0.46$, $0.49$] 0.76 [$0.75$, $0.77$]
Extrusion F1 $\uparrow$ 0.98 [$0.98$, $0.98$] 0.90 [$0.89$, $0.90$] 0.81 [$0.80$, $0.81$] 0.94 [$0.93$, $0.94$]
CD mean $\downarrow$ 50.44 [$46.99$, $53.92$] 144.85 [$140.06$, $149.69$] 107.47 [$103.45$, $111.48$] 28.52 [$26.58$, $30.47$]
CD median $\downarrow$ 0.31 [$0.29$, $0.32$] 50.78 [$46.96$, $55.46$] 33.38 [$30.78$, $35.88$] 0.37 [$0.34$, $0.40$]
SIR $\downarrow$ 0.04 [$0.04$, $0.05$] 0.04 [$0.04$, $0.05$] 0.08 [$0.07$, $0.08$] 0.04 [$0.04$, $0.05$]
DangEL $\downarrow$ 1.36 [$1.22$, $1.51$] 1.11 [$0.97$, $1.26$] 2.38 [$2.17$, $2.60$] 1.06 [$0.94$, $1.19$]
SegE $\downarrow$ 0.63 [$0.53$, $0.78$] 0.50 [$0.44$, $0.57$] 1.06 [$0.97$, $1.16$] 0.48 [$0.44$, $0.53$]
FluxEE ($\times 10^{2}$) $\downarrow$ 0.39 [$0.31$, $0.47$] 0.57 [$0.45$, $0.71$] 0.93 [$0.80$, $1.06$] 0.42 [$0.33$, $0.53$]
Watertightness (%) $\uparrow$ 87.63 [$86.88$, $88.36$] 91.90 [$91.29$, $92.50$] 82.48 [$81.62$, $83.34$] 90.52 [$89.85$, $91.17$]
@@
EECM $\uparrow$ 0.87 [$0.86$, $0.88$] 0.77 [$0.78$, $0.80$] 0.66 [$0.66$, $0.68$] 0.82 [$0.83$, $0.85$]
DMCD ($\times 10^{3}$) $\downarrow$ 3.37 [$3.16$, $3.59$] 7.64 [$7.32$, $7.95$] 9.19 [$8.89$, $9.50$] 5.09 [$4.81$, $5.36$]
SD mean ($\times 10^{2}$) $\downarrow$ 2.50 [$2.34$, $2.68$] 8.31 [$8.01$, $8.62$] 11.31 [$10.91$, $11.71$] 4.68 [$4.42$, $4.95$]
SD median ($\times 10^{2}$) $\downarrow$ 0.00 [$0.00$, $0.00$] 3.20 [$2.87$, $3.48$] 5.07 [$4.72$, $5.48$] 0.02 [$0.00$, $0.05$]